The work done by lifting a body of mass 20kg and specific gravity 3.2 to a height of 8m in water is?(g=10ms-2)

Grade:12

1 Answers

Srivathsa
5 years ago
Bouyant force =Mg/density =200/3.2=62.5
Weight of body in water=Weight - bouyant force=200-62.5=137.5
 
W=F•S=FScosx=137.5×8×1=1100J
✓✓